A box contains 4 red, 4 white and 5 green balls.Three balls are drawn from the
box together. Find the probability that they may be
i) all the same color
ii) all different colors?

Answer by Edwin McCravy(20077)

Find the probability that they may be
i) all of same color.
There are 4+4+5=13 balls.  So the number of ways to draw any 3 of them is
13C3 = (13∙12∙11)/(3∙2∙1) = 286 ways.

The number of ways they can be all red is 4C3 = (4∙3∙2)/(3∙2∙1) = 4 ways.

The number of ways they can be all white is 4C3 = (4∙3∙2)/(3∙2∙1) = 4 ways.

The number of ways they can be all green is 5C3 = (5∙4∙3)/(3∙2∙1) = 10 ways.

That's 4+4+10=18 ways they call all be the same color.

That's 18 ways out of 286 or a probability of 18/286 = 9/143

-------------------------------------------------

Find the probability that they may be
ii) all of different colors?

Choose the red ball any of 4 ways.
Choose the white ball any of 4 ways.
Choose the green ball any of 5 ways.

That's 4∙4∙5 = 80 ways to choose the balls all different colors.

That's 80 ways out of 286 or a probability of 80/286 = 40/143

Edwin